# Does AI Code Review Lead to Code Changes? A Case Study of GitHub Actions

### v1

@@ Abstract
AI-based code review tools automatically review and comment on pull requests to improve code quality. Despite their growing presence, little is known about their actual impact. We present a large-scale empirical study of 16 popular AI-based code review actions for GitHub workflows, analyzing more than 22,000 review comments in 178 repositories. We investigate (1) how these tools are adopted and configured, (2) whether their comments lead to code changes, and (3) which factors influence their effectiveness. We develop a two-stage LLM-assisted framework to determine whether review comments are addressed, and use interpretable machine learning to identify influencing factors. Our findings show that, while adoption is growing, effectiveness varies widely. Comments that are concise, contain code snippets, and are manually triggered, particularly those from hunk-level review tools, are more likely to result in code changes. These results highlight the importance of careful tool design and suggest directions for improving AI-based code review systems.
Index Terms:
Code Review, GitHub Actions, Large Language Models, Empirical Software Engineering
@@ I Introduction
Code review is a widely adopted practice in modern software engineering, playing a central role in identifying bugs, improving code quality, and facilitating knowledge transfer within teams [1]. Traditionally, code reviews are performed manually, with developers reviewing and commenting on code changes submitted via pull requests [2]. However, recent advances in generative AI (GenAI) have created opportunities to automate parts of this process by generating human-like code review comments [3].
On platforms such as GitHub, developers can now integrate GenAI into their workflows via GitHub Actions [4]. These AI-driven code review actions automatically analyse code changes and post suggestions as comments on pull requests. Several of such actions are available on the GitHub Marketplace and have gained popularity, promising to reduce developer effort and to increase review coverage [5]. Despite their growing presence, there is little empirical evidence of how these tools are actually used in practice and whether they meaningfully impact the development process.
Previous research has examined human code review practices in detail [6], including factors that make review comments useful [7, 8, 9], the granularity of feedback [10], and developer responsiveness [11]. In contrast, the dynamics of automated reviews powered by GenAI remain largely unexplored. Open questions include: Which AI-based tools are developers actually adopting? Do generated comments lead to meaningful code changes? Which characteristics influence whether such comments are taken seriously and acted upon?
This paper addresses these gaps through an empirical study of AI-based code review actions based on Large Language Models (LLMs) on GitHub. Our goal is to understand both the usage and the effectiveness of these tools, as well as the factors that influence their impact. Specifically, we pose the following research questions:
RQ1
How are LLM-driven code review actions adopted in GitHub repositories?
@@
RQ2
How accurately can we automatically detect whether code review comments have been addressed?
Our second question addresses the challenge of measuring the effectiveness of AI-generated comments. Manually assessing whether review suggestions are implemented at scale is infeasible. We therefore develop and evaluate an automated method to detect whether comments have been addressed, providing a scalable approach to quantify comment impact.
RQ3
@@
Finally, we explore the contextual and technical factors that make comments more or less likely to result in code changes. By analysing aspects such as comment content, contributor experience, and project characteristics, we aim at identifying patterns that can inform the design of more effective review tools.
@@
To answer these questions, we conducted a large-scale empirical study of 16 popular GenAI-based code review actions on GitHub. We analyse their adoption in 178 repositories, collect more than 22,000 review comments, and develop a two-stage LLM-based classification framework to assess whether these comments are addressed. Finally, we use interpretable machine learning techniques to model the factors that influence the effectiveness of comments.
@@
  • We provide the first systematic study of the adoption and usage of AI-based code review actions on GitHub, categorising tools by their review granularity (i.e., PR, file, or hunk) and behaviour (e.g., action triggers).
@@ III Study Setup
To investigate AI-based code review on GitHub, we collected the 16 most popular actions as our study sample and conducted a categorization based on their review behaviors. We provide an online appendix [12] (to be published on Zenodo after acceptance), including our dataset, annotations, and scripts for LLM-assisted analysis of whether review comments have been addressed (for RQ2), and for the interpretation of how factors impact the likelihood of comment addressing (for RQ3). We follow the guidelines proposed by Wagner et al. [13, 14] for empirical studies in software engineering involving LLMs.
@@ IV Analyses & Results
IV-A RQ1: How are LLM-driven code review actions adopted in GitHub repositories?
@@
TABLE IV: RQ1: Adoption statistics of 16 AI-based review actions, grouped by PR/File/Hunk review actions. Total/mature refers to activation of action, the last column lists review activity. Asterisk (*) indicates general instead of inline comments.
ID Total Repos Mature Repos Observed Review Activity (Number of Repos, PRs, Comments)
@@
5 7 2 Repos=0, PRs=0, Comments=0
6 4 1 Repos=0, PRs=0, Comments=0
8 3 1 Repos=1, PRs=24, Comments*=37
9 5 1 Repos=0, PRs=0, Comments=0
11 8 4 Repos=1, PRs=38, Comments*=38
16 13 5 Repos=5, PRs=74, Comments*=133
Total 40 14 Repos=7, PRs=136, Comments=208
@@
1 460 114 Repos=74, PRs=2,238, Comments=19,549
2 37 11 Repos=9, PRs=421, Comments*=421
13 3 1 Repos=0, PRs=0, Comments=0
15 4 1 Repos=1, PRs=1, Comments=3
Total 504 127 Repos=84, PRs=2,660, Comments=19,973
@@
3 108 27 Repos=14, PRs=171, Comments=1,911
4 47 8 Repos=5, PRs=28, Comments=193
7 5 0 Repos=0, PRs=0, Comments=0
10 6 1 Repos=1, PRs=4, Comments=31
12 6 1 Repos=1, PRs=3, Comments=10
14 2 0 Repos=0, PRs=0, Comments=0
Total 174 37 Repos=21, PRs=206, Comments=2,145
All 16 718 178 Repos=112, PRs=3,002, Comments=22,326
@@
IV-B RQ2: How accurately can we automatically detect whether code review comments have been addressed?
Approach: To assess the practical impact of these actions, we leverage LLMs to extensively examine whether the generated review comments led to actual code changes (which we define as “addressed”). We use the term “addressing analysis” to refer to our analysis of whether comments have been addressed and we use the term “addressing dataset” to refer to the dataset we created (filtered from RQ1 and augmented with human-written review comments) to carry out this analysis. To establish ground truth for this addressing dataset, we manually annotated a sampled subset and evaluated several LLMs on the annotated data to identify the most effective approach for automating this assessment at scale.
Phase I: Refining subset for Addressing Analysis. We first excluded comments from PR-level actions because of their significant differences from typical human code review practices and their low volume. For comments from file-level and hunk-level actions, we only retained those in merged PRs (and prior to merging) to capture finalized developer decisions, yielding 16,762 comments. To facilitate human analysis, we selected only English comments using FastText, reducing the dataset to 4,229 comments (a removal of 12,533 comments). ID-1 (anc95/ChatGPT-CodeReview) saw a significant reduction in comments at this stage, with 11,744 comments (mostly in Korean) filtered out. We further discuss potential threats introduced by this language-based filtering in Section VI. Given that some actions enable user-LLM dialogue within threads, we restricted our analysis to first-in-thread comments to focus on the original review suggestions rather than follow-up interactions (excluding an additional 34 comments). After these filters, only Actions ID-1 to ID-4 retained sufficient data for the addressing analysis, with 2,908, 399, 713, and 169 comments respectively.
@@ IV-B RQ2: How accurately can we automatically detect whether code review comments have been addressed?
TABLE VI: RQ2: Annotation Label Distribution by Comment Sources.
Phase II: Adding Human-Authored Review Comments for Comparison. For comparison, we collected human-authored review comments from the same 51 repositories for which we had obtained action-generated comments in Phase I. For each repository, we defined its AI review activation period as the time span between the first and last action-generated comment. We restricted our collection of human-written comments to this same time window to minimize potential effects on the addressing ratio (e.g., differences in the project’s development phase). Within this period, we collected human-authored inline comments that met the same filtering criteria as the action-generated comments: (1) from merged pull requests and posted prior to merging, (2) written in English, and (3) first in their comment threads (this filter excluded nearly half of the candidate comments). We also excluded comments from accounts whose login names contain the string “bot” [20]. This process yielded a final set of 1,166 human-authored review comments. Following the same procedure used for hunk-level review actions, we reconstructed their originally reviewed changes and subsequent file modifications. Table V summarizes the statistics for the human comments. In total, our dataset for addressing analysis includes 5,652 review comments: 3,604 from file-level review actions, 882 from hunk-level actions, and 1,166 authored by human reviewers.
Phase III: Human Annotation. To assess whether code review comments lead to code changes, we randomly sampled 50 comments from each category where subsequent file modifications occurred (i.e. 467 eligible from file-level actions, 424 eligible from hunk-level actions, and 1,070 eligible human-authored comments), yielding 150 comments in total. Before the full annotation, two authors first established an annotation guideline through consensus on 9 comments (the first 3 from each category). Through this process, we identified two challenges that prevent comment addressing from being a simple binary classification. First, some comments did not contain any suggestions, such as simple praise in human reviews (e.g., “Amazing!”), or mere code summaries in action-generated reviews. Some comments contained overly general suggestions lacking clear guidance (e.g., “Suggest to test thoroughly”), making it difficult to determine whether they were addressed. Second, a single comment may include multiple issues or suggestions, requiring a more granular assessment of their resolution. Therefore, we designed a two-stage annotation task. In Stage-1, we examined the comment text and the reviewed changes to determine whether a comment contained valid (i.e. specific and actionable) issues or suggestions, resulting in the following classifications:
@@
Following this guideline, the remaining 141 comments (47 per category) were independently labeled by one author and an external graduate student in software engineering who is not a co-author of this paper. Inter-rater agreement (Cohen’s $\kappa$) reached substantial levels over the full 6-class annotation scheme (i.e., “None”, “General”, “Valid-Uncertain”, “Valid-Unaddressed”, “Valid-Partially”, “Valid-Fully”): 0.674 for file-level actions, 0.734 for hunk-level actions, and 0.764 for human reviews. A third author resolved disagreements by selecting between two annotations or making a new judgment when necessary. Table VI shows the final label distribution across 150 annotated comments.
@@
Phase IV: LLM Evaluation for Addressing Analysis. To automate the detection of comment addressing, we evaluated three families of LLMs on our human-annotated dataset: OpenAI [16] (gpt-4.1, gpt-4o, and o4-mini / o3-mini with medium reasoning effort), Claude [18] (claude-3-sonnet and claude-3-haiku), and DeepSeek [21] (deepseek-r1 and deepseek-v3). The specific API endpoints and model versions used are documented in our online appendix scripts; the evaluations were performed in April and May 2025. Following the human annotation process, we designed separate prompts for Stage-1 and Stage-2. Stage-1 aimed to classify comments as “None”, “General”, or “Valid”, and extract a list of valid issues or suggestions for “Valid” comments. When Stage-1 identified a comment as “Valid”, Stage-2 was triggered to determine its final addressed status. Both stages utilized the same contextual information (i.e. reviewed changes and subsequent file modifications) and the classification scheme as the human annotators. The details of the LLM-assisted framework with specific prompts are available in the online appendix for other researchers to use.
@@
TABLE VII: RQ2: Performance on 150 annotated comments using the best automated setup, with openai-gpt-4.1 for Stage-1 and openai-o3-mini for Stage-2. “Avg.”: average across three sources.
@@
Source Full 6-class Stage-1 Stage-2 OA Cohen’s $\kappa$ OA Cohen’s $\kappa$ OA Cohen’s $\kappa$ File-level Action 83.6% 76.4% 94.4% 86.9% 88.5% 77.1% Hunk-level Action 89.2% 84.7% 95.2% 81.7% 93.2% 86.3% Human 85.6% 68.9% 94.0% 76.2% 93.7% 68.2% Avg. 86.1% 76.7% 94.5% 81.6% 91.8% 77.2%
@@
We decoupled the stages to allow different model selection based on stage-specific performance. With temperature set to 0, we ran each model five times for robust evaluation. Performance was measured using overall accuracy (OA) and Cohen’s $\kappa$. In the first round of evaluation, we used the same LLM for both Stage 1 and Stage 2. Beyond assessing performance on the final 6-class scheme, we also evaluated the models’ effectiveness in each stage independently. For Stage-1 (detecting valid comment), we grouped “None” and “General” into a “Not Valid” category, and reported the models’ ability to distinguish between “Not Valid” and “Valid” comments. For Stage-2 (addressed assessment), focusing on comments correctly identified as “Valid”, we grouped “Valid-Uncertain”, “Valid-Unaddressed” into “Valid, Not Addressed,” and “Valid-Partially” and “Valid-Fully” into “Valid, Addressed”, then reported the models’ ability to differentiate these two outcomes. In the second round of evaluation, we cross-combined the top three models from each stage to find the optimal combination.
Results: In our first round of evaluation, we observed that general-purpose models usually performed well in Stage-1, while reasoning-enhanced models excelled in Stage-2. The top performers for Stage-1 were gpt-4.1 (94.5% average overall accuracy), deepseek-v3 (94.0%), and claude-3-sonnet (93.2%). For Stage-2, deepseek-r1 led with 95.4%, followed by openai-o3-mini (92.4%) and o4-mini (91.3%). Table VII reports the performance of our optimal cross-combined setup: gpt-4.1 for Stage-1 and o3-mini for Stage-2, which achieved strong and balanced performances across three comment sources, with an average of 86.1% overall accuracy and 76.7% Cohen’s $\kappa$ under the full 6-class scheme. We observe that the human-authored comments show a relatively lower Cohen’s $\kappa$ (e.g. 68.9% for full 6-class classification). This can be attributed to the highly imbalanced label distribution in the human-authored comments, where 80% fall into the “Valid-Fully” category (see Table VI). In such imbalanced scenarios, $\kappa$ statistics become more sensitive to disagreements. Nevertheless, our automated approach achieves substantial agreement levels across all three comment categories, with Cohen’s $\kappa$ ranging from 68.9% to 84.7%, demonstrating that we can effectively assess comment validity and addressing status for subsequent analysis.
Answer to RQ2: Our best-performing configuration achieved 94.5% accuracy in identifying valid comments (Stage-1), 91.8% accuracy in assessing addressing status (Stage-2), and a strong overall accuracy of 86.1% across the full 6-class classification task.
@@
TABLE VIII: RQ3: LLM-Assigned Addressing Labels for Comments by Source (ID-1: anc95/ChatGPT-CodeReview, ID-2: mattzcarey /code-review-gpt, ID-3: coderabbitai/ai-pr-reviewer, ID-4: aidar-freeed/ai-codereviewer, Human Reviews). Distribution shift compared with Table VI results from a different sampling strategy (see RQ3’s results).

### v2

@@ Abstract
AI-based code-review tools automatically review and comment on pull requests to improve code quality. Despite their growing presence, little is known about their actual impact. We present a large-scale empirical study of 16 popular AI-based code-review actions for GitHub workflows, analyzing more than 22,000 review comments in 178 repositories. We investigate (1) how these tools are adopted and configured, (2) whether their comments lead to code changes, and (3) which factors influence their effectiveness. We develop a two-stage LLM-assisted framework to determine whether review comments are addressed. We then use interpretable machine learning to identify the influencing factors. We found that while adoption is growing, its effectiveness varies widely. Comments that are concise, contain code snippets, and are manually triggered, particularly those from hunk-level review tools, are more likely to result in code changes. These results highlight the importance of tool design and suggest directions for improving AI-based code review systems.
@@ I Introduction
Code review is a widely adopted practice in modern software engineering, playing a central role in identifying bugs, improving code quality, and facilitating knowledge transfer within teams [42]. Traditionally, code reviews are performed manually through reviewing and commenting on code changes (submitted via pull requests) by developers [19]. However, recent advances in generative AI (GenAI) have created opportunities to automate at least part of this process by generating human-like code review comments [33].
On platforms such as GitHub, developers can now integrate GenAI into their workflows via GitHub Actions [51]. These AI-driven code review actions automatically analyse code changes and post suggestions as comments on pull requests. Some of these actions are now available and popular on the GitHub Marketplace, promising to reduce developers’ efforts and to increase review coverage [9]. Despite the growing presence, there is little empirical evidence of how these actions are actually used in practice and whether they can impact the development process meaningfully.
Previous research has investigated human code review practices in detail [2], including (1) factors that make review comments useful [6, 47, 39], (2) the granularity of feedback [31], and (3) developer responsiveness [35]. In contrast, the dynamics of automated reviews powered by GenAI remain largely unexplored. Open questions include: Which AI-based tools are developers actually adopting? Do generated comments lead to meaningful code changes? Which characteristics influence whether such comments are taken seriously and acted upon?
This paper addresses these gaps through an empirical study of AI-based code review actions based on Large Language Models (LLMs) on GitHub. Our goal is to understand both the usage and the effectiveness of these tools, as well as the factors that influence their impact. Specifically, we propose the following three research questions:
RQ1:
How are LLM-based code review actions adopted in GitHub repositories?
@@
RQ2:
To what extent do AI-generated review comments lead to code changes compared to human review comments?
Our second question aims to understand whether AI-generated review comments can truly impact the code (i.e., lead to code changes) and how their effectiveness compares to human review comments. To enable large-scale analysis, we propose an LLM-based approach to automatically assess the validity of review comments and whether they had been addressed.
RQ3:
@@
Finally, we explore the contextual and technical factors that make comments more or less likely to result in code changes. By analyzing aspects such as comment content, contributor experience, and project characteristics, we aim at identifying patterns that can inform the design of more effective review tools.
@@
To answer these questions, we conducted a large-scale empirical study of 16 popular GenAI-based code-review actions on GitHub. We analyze their adoption in 178 repositories, collect more than 22,000 review comments, and develop a two-stage LLM-based classification framework to assess whether these comments are addressed. Finally, we use interpretable machine learning techniques to model and find out the factors that influence the effectiveness of comments.
@@
  • We provide the first systematic study of the adoption and usage of AI-based code review actions on GitHub, categorizing tools by their review granularity (i.e., PR, file, or hunk) and behaviour (e.g., action triggers).
@@ III Study Setup
To investigate AI-based code review on GitHub, we collected the 16 most popular actions as our study sample and conducted a categorization based on their review behaviors. We provide an online appendix [13], including our dataset, annotations, and scripts for LLM-assisted analysis of whether review comments have been addressed (for RQ2), and for the interpretation of how factors impact the likelihood of comment addressing (for RQ3). We follow the guidelines proposed by Wagner et al. [49, 3] for empirical studies in software engineering involving LLMs.
@@ IV Analyses & Results
IV-A RQ1: How are LLM-based code review actions adopted in GitHub repositories?
@@
TABLE IV: RQ1: Adoption statistics of 16 AI-based review actions, grouped by PR/File/Hunk review actions. “Total/Mature Repos” counts repositories configuring the action (Mature: $\geq$50 PRs). The observed activity columns detail the adoption distribution across repos, PRs, and comments in mature repos. Asterisk (*) indicates general instead of inline comments.
ID Total Repos Mature Repos Observed Review Activity
Repos PRs Comments
@@
5 7 2 0 0 0
6 4 1 0 0 0
8 3 1 1 24 37*
9 5 1 0 0 0
11 8 4 1 38 38*
16 13 5 5 74 133*
Total 40 14 7 136 208
@@
1 460 114 74 2,238 19,549
2 37 11 9 421 421*
13 3 1 0 0 0
15 4 1 1 1 3
Total 504 127 84 2,660 19,973
@@
3 108 27 14 171 1,911
4 47 8 5 28 193
7 5 0 0 0 0
10 6 1 1 4 31
12 6 1 1 3 10
14 2 0 0 0 0
Total 174 37 21 206 2,145
All 16 718 178 112 3,002 22,326
@@
IV-B RQ2: To what extent do AI-generated review comments lead to code changes compared to human review comments?
Approach: To assess the practical impact of these actions, we leverage LLMs to extensively examine whether the generated review comments led to actual code changes (which we define as “addressed”). We use the term “comment addressing analysis” to refer to our analysis of whether comments have been addressed. We first created the “comment addressing dataset” (filtered from RQ1 and augmented with human-written review comments) to carry out this analysis. Next, we sampled a subset to manually annotate whether they had been addressed and evaluated LLM performance on this subset. Finally, we employed the best-performing LLM approach across the entire dataset. Figure 3 presents an overview of this workflow.
Figure 3: Overview of the RQ2 comment addressing analysis pipeline.
Phase I: Refining the Subset for Comment Addressing Analysis. We first excluded comments from PR-level actions because of their significant differences from typical human code review practices and their low volume. For comments from file-level and hunk-level actions, we only retained those in merged PRs (and prior to merging) to capture finalized developer decisions, yielding 16,816 comments. To facilitate human analysis, we selected only English comments using FastText, reducing the dataset to 4,229 comments (a removal of 12,587 comments). ID-1 (anc95/ChatGPT-CodeReview) saw a significant reduction in comments at this stage, with 11,744 comments (mostly in Korean) filtered out. We further discuss potential threats introduced by this language-based filtering in Section VI. Given that some actions enable user-LLM dialogue within threads, we restricted our analysis to first-in-thread comments to focus on the original review suggestions rather than follow-up interactions (excluding an additional 34 comments). After these filters, only Actions ID-1 to ID-4 retained sufficient data for the addressing analysis, with 2,908, 399, 713, and 169 comments respectively.
@@
For comparison, we collected human-authored review comments from the same 51 repositories for which we had obtained action-generated comments in Phase I. For each repository, we defined its AI review activation period as the time span between the first and last action-generated comment. We restricted our collection of human-written comments to this same time window to minimize potential effects on the addressing ratio (e.g., differences in the project’s development phase). Within this period, we collected human-authored inline comments that met the same filtering criteria as the action-generated comments: (1) from merged pull requests and posted prior to merging, (2) written in English, and (3) first in their comment threads (this filter excluded nearly half of the candidate comments). We also excluded comments from accounts whose login names contain the string “bot” [18]. This process yielded a final set of 1,166 human-authored review comments. Following the same procedure used for hunk-level review actions, we reconstructed their originally reviewed changes and subsequent file modifications. Table V summarizes the statistics for the human comments. In total, our comment addressing dataset includes 5,652 review comments: 3,604 from file-level review actions, 882 from hunk-level actions, and 1,166 authored by human reviewers.
Phase II: Establishing Ground Truth via Human Annotation. In this phase, we sampled a subset and manually evaluated whether they had been addressed to serve as the ground truth. Specifically, we randomly sampled 50 comments from each category where subsequent file modifications occurred (i.e. 467 eligible from file-level actions, 424 eligible from hunk-level actions, and 1,070 eligible human-authored comments), yielding 150 comments in total. Before the full annotation, two authors first established an annotation guideline through consensus on 9 comments (the first 3 from each category). Through this process, we identified two challenges that prevent comment addressing from being a simple binary classification. First, some comments did not contain any suggestions, such as simple praise in human reviews (e.g., “Amazing!”), or mere code summaries in action-generated reviews. Some comments contained overly general suggestions lacking clear guidance (e.g., “Suggest to test thoroughly”), making it difficult to determine whether they were addressed. Second, a single comment may include multiple issues or suggestions, requiring a more granular assessment of their resolution. Therefore, we designed a two-stage annotation task. In Stage-1, we examined the comment text and the reviewed changes to determine whether a comment contained valid (i.e. specific and actionable) issues or suggestions, resulting in the following classifications:
@@
Following this guideline, the remaining 141 comments (47 per category) were independently labeled by one author and an external graduate student in software engineering who is not a co-author of this paper. Inter-rater agreement (Cohen’s $\kappa$) reached substantial levels over the full 6-class annotation scheme (i.e., “None”, “General”, “Valid-Uncertain”, “Valid-Unaddressed”, “Valid-Partially”, “Valid-Fully”): 0.674 for file-level actions, 0.734 for hunk-level actions, and 0.764 for human reviews. A third author resolved disagreements by selecting between two annotations or making a new judgment when necessary. Table VIII shows the final label distribution across 150 annotated comments.
@@
Phase III: Evaluating LLM Effectiveness for Comment Addressing Analysis. Based on our human-annotated dataset, we evaluated the performance of three families of LLMs for addressing analysis: OpenAI [36] (gpt-4.1, gpt-4o, and o4-mini / o3-mini with medium reasoning effort), Claude [8] (claude-3-sonnet and claude-3-haiku), and DeepSeek [10] (deepseek-r1 and deepseek-v3). The specific API endpoints and model versions used are documented in our online appendix scripts; the evaluations were performed in April and May 2025. Following the human annotation process, we designed separate prompts for Stage-1 and Stage-2. Stage-1 aimed to classify comments as “None”, “General”, or “Valid”, and extract a list of valid issues or suggestions for “Valid” comments. When Stage-1 identified a comment as “Valid”, Stage-2 was triggered to determine its final addressed status. Both stages utilized the same contextual information (i.e. reviewed changes and subsequent file modifications) and the classification scheme as the human annotators. The details of the LLM-assisted framework with specific prompts are available in the online appendix for other researchers to use.
TABLE VI: RQ2: Annotation Label Distribution by Comment Sources.
@@
TABLE VII: RQ2: Performance on 150 annotated comments using the best automated setup, with openai-gpt-4.1 for Stage-1 and openai-o3-mini for Stage-2. The last column reports the ability to distinguish “Valid, Addressed” comments from all others. “Avg.”: average across three sources.
@@
Source Full 6-class Stage-1 (Validity) Stage-2 (Addressing) Valid, Addressed vs. All Others OA Macro-F1 OA Macro-F1 OA Macro-F1 OA Macro-F1 File-level Action 83.6% 78.1% 94.4% 93.5% 88.5% 88.4% 92.0% 90.5% Hunk-level Action 89.2% 83.4% 95.2% 90.8% 93.2% 93.1% 91.6% 91.2% Human 85.6% 62.4% 94.0% 88.0% 93.7% 83.9% 90.8% 87.6% Avg. 86.1% 74.6% 94.5% 90.8% 91.8% 88.5% 91.5% 89.8%
TABLE VIII: RQ2: Performance of distinguishing “Valid, Addressed” comments from others across the entire dataset. Unchanged Group is logically classified as non-addressed (100% Acc.).
Source No Code Change With Code Change Overall
Total Acc. Total Acc. Total Acc.
File-level Action 3,137 100% 467 92.0% 3,604 99.0%
Hunk-level Action 458 100% 424 91.6% 882 96.0%
Human 96 100% 1,070 90.8% 1,166 91.6%
Total 3,691 100% 1,961 91.3% 5,652 97.0%
@@
We decoupled the stages to allow different model selection based on stage-specific performance. With temperature set to 0, we ran each model five times for robust evaluation. Performance was measured using overall accuracy (OA) and Macro-F1 scores. In the first round of evaluation, we used the same LLM for both Stage 1 and Stage 2. Beyond assessing performance on the final 6-class scheme, we also evaluated the models’ effectiveness in each stage independently. For Stage-1 (detecting valid comment), we grouped “None” and “General” into a “Not Valid” category, and reported the models’ ability to distinguish between “Not Valid” and “Valid” comments. For Stage-2 (addressed assessment), focusing on comments correctly identified as “Valid”, we grouped “Valid-Uncertain”, “Valid-Unaddressed” into “Valid, Not Addressed,” and “Valid-Partially” and “Valid-Fully” into “Valid, Addressed”, then reported the models’ ability to differentiate these two outcomes. In the second round of evaluation, we cross-combined the top three models from each stage to find the optimal combination.
@@
TABLE IX: RQ2: LLM-Assigned Addressing Labels for Comments by Source (ID-1: anc95/ChatGPT-CodeReview, ID-2: mattzcarey /code-review-gpt, ID-3: coderabbitai/ai-pr-reviewer, ID-4: aidar-freeed/ai-codereviewer, Human Reviews). Distribution shift compared with Table VIII results from a different sampling strategy (see RQ2’s results).
@@
Phase IV: LLM-Based Labeling of Comment Addressing. Finally, we apply the best-performing setup determined in Phase III to conduct comment addressing analysis on the entire dataset of 5,652 review comments.
Results: We first present the evaluation results from Phase III. In our first round of evaluation, we observed that general-purpose models usually performed well in Stage-1, while reasoning-enhanced models excelled in Stage-2. The top performers for Stage-1 were gpt-4.1 (94.5% average overall accuracy), deepseek-v3 (94.0%), and claude-3-sonnet (93.2%). For Stage-2, deepseek-r1 led with 95.4%, followed by openai-o3-mini (92.4%) and o4-mini (91.3%). Table VIII reports the performance of our optimal cross-combined setup: gpt-4.1 for Stage-1 and o3-mini for Stage-2, which achieved strong and balanced performances across three comment sources, with an average of 86.1% overall accuracy and 74.6% Macro-F1 score under the full 6-class scheme. The lower macro-F1 for human-authored comments (62.4%) reflects stronger class imbalance within that subset (see Table VIII), which penalizes macro-averaged metrics. Importantly, performance on the primary downstream distinction (“Valid, Addressed” vs. all others) remains high, with an average of 91.5% overall accuracy and 89.8% macro-F1. Furthermore, when analyzing the complete dataset, we found that a significant portion of comments (87.0% of file-level, 51.9% of hunk-level, and 8.2% of human comments) had no subsequent modifications. We can thus trivially classify them as non-“Valid, Addressed” with certainty. Considering the natural class imbalance in real-world review data, the overall classification accuracy across the full dataset is 97%, ensuring the reliability of our subsequent large-scale analysis in Phase IV (detailed breakdown in Table VIII).
Following the best-performing setup on the annotated dataset, we used gpt-4.1 to assess the validity of all 5,652 comments, identifying 3,955 as valid, and then used openai-o3-mini to assess whether these valid comments had been addressed. Table IX details the results across the entire dataset. Comparing Table VIII (manual labels) and Table IX (automated labels), we observe a discrepancy between their annotation distributions. This stems from our sampling strategy. For the manual annotation, we purposefully sampled only those comments where the reviewed file was subsequently modified. This was a necessary precondition for assessing if a comment was addressed, but it naturally biased the sample toward “Valid” and “Addressed” classifications. In contrast, the automated annotation was applied to the entire comment dataset, which included a large volume of comments (65.3%) with no subsequent file changes. Following our procedure, if these comments were identified as “Valid” in Stage-1, they were automatically classified as “Not Addressed” in Stage-2. Consequently, the proportion of “Valid, Addressed” comments is lower in the automated setting. Beyond this discrepancy, both tables indicate that the addressing rate of AI-generated review comments (0.9%–19.2% in Table IX) still lags behind human review comments (60%). Overall, hunk-level review actions exhibit a higher addressing rate (6.5%–19.2%) compared to file-level actions (0.9%–4.2%).
Answer to RQ2: We found that hunk-level review actions (6.5%–19.2%) exhibit a higher addressing rate compared to file-level actions (0.9%–4.2%), yet the addressing rate of AI-generated review comments still lags behind human review comments (60%).